### v1

# Leveraging AI Agents for Autonomous Networks: A Reference Architecture and Empirical Studies

Abstract
The evolution toward Level 4 (L4) Autonomous Networks (AN) represents a strategic inflection point in telecommunications, where networks must transcend reactive automation to achieve genuine cognitive capabilities—fulfilling TM Forum’s vision of self-configuring, self-healing, and self-optimizing systems that deliver zero-wait, zero-touch, and zero-fault services. This work bridges the gap between architectural theory and operational reality by implementing Joseph Sifakis’s AN Agent reference architecture in a functional cognitive system, deploying coordinated proactive-reactive runtimes driven by hybrid knowledge representation. Through an empirical case study of a Radio Access Network (RAN) Link Adaptation (LA) Agent, we validate this framework’s transformative potential: demonstrating sub-10 ms real-time control in 5G NR sub-6 GHz while achieving 6% higher downlink throughput than Outer Loop Link Adaptation (OLLA) algorithms and 67% Block Error Rate (BLER) reduction for ultra-reliable services through dynamic Modulation and Coding Scheme (MCS) optimization. These improvements confirm the architecture’s viability in overcoming traditional autonomy barriers and advancing critical L4-enabling capabilities toward next-generation objectives.
publicationid: pubid:
I Introduction
Autonomous Networks (AN), a purpose-specific telecommunications technology pioneered by the TM Forum (TMF) in 2019, target networks with intrinsic self-configuration, self-healing, and self-optimization capabilities—collectively termed the Three-Self Capabilities [1]. These fundamental properties enable the realization of zero-wait, zero-touch, and zero-fault network services, known as the Three-Zero Objectives, which collectively deliver optimal user experiences while maximizing resource utilization throughout the entire network lifecycle. By strategically integrating emerging general-purpose technologies including artificial intelligence (AI), digital twins, and big data analytics, AN not only transforms conventional network operations but fundamentally reorients value creation paradigms from traditional device-centric and management-centric models toward customer-oriented, service-driven, and business-focused frameworks.
The TMF’s standardized six-level maturity continuum (L0–L5) provides telecommunications operators with a structured methodology for assessing AN evolutionary progress, where ascending levels denote progressively enhanced automation and cognitive capabilities. Currently, the global telecommunications industry is undergoing a critical transition phase, advancing from mid-to-low autonomy stages (L2/L3: partial or conditional autonomy) toward the pivotal threshold of L4 (high-level autonomy). Major operators worldwide have established comprehensive roadmaps aligning with this evolutionary trajectory. Nevertheless, AN implementation confronts significant systemic challenges as a complex systems engineering endeavor. Persistent limitations in AI algorithmic complexity and technological maturity have engendered operational bottlenecks metaphorically described as “intelligence plateau” and “acceleration resistance.” Traditional machine learning-based approaches embed task-specific intelligence into AN workflows but exhibit inherent deficiencies in genuine autonomous decision-making. Concurrently, Large Language Model (LLM)-based AI copilots demonstrate enhanced knowledge provision and recommendation capabilities through human-AI collaboration frameworks, yet remain constrained by limited proactive intervention capacities. Fundamentally, both methodologies encounter autonomy ceilings at conditional levels (L3), necessitating the development of agents with advanced self-governance mechanisms to achieve L4 operational targets.
The conceptual foundation for autonomous agents traces back to Marvin Minsky’s seminal 1986 work The Society of Mind [2], which introduced the notion of computational entities exhibiting intelligent behavior. Contemporary autonomous agents represent sophisticated computational systems capable of demonstrating contextually appropriate behaviors within specific operational environments. These systems integrate core competencies spanning environmental perception, cognitive reasoning, strategic decision-making, and physical actuation, continuously enhancing performance through experiential learning and adaptive recalibration. By leveraging accumulated knowledge repositories, engaging in self-directed knowledge acquisition, and participating in collaborative multi-agent ecosystems, they autonomously accomplish predefined operational objectives.
While reinforcement learning (RL) agents previously dominated autonomous systems—demonstrating adaptive behaviors through iterative environmental interactions yet fundamentally limited by their reactive nature and lack of higher-order cognition—the emergence of generative AI has catalyzed a paradigm shift toward LLM-based agents, marking a significant evolutionary leap toward genuinely autonomous and cognitively sophisticated AI systems. Modern LLM-based agents exhibit proactive cognitive engagement, requiring minimal supervisory intervention while integrating multifaceted capabilities including knowledge synthesis, associative memory, logical reasoning, multi-step planning, and risk-aware decision-making. These systems actively utilize digital tools and environmental interfaces for contextual judgment and execution, with continuous learning mechanisms enabling knowledge expansion, adaptation to novel operational scenarios, and pursuit of emergent objectives—collectively establishing the technical underpinnings for high-level operational autonomy.
This research contributes two principal advancements to the field: first, it presents a novel dual-driver AN Agent reference architecture that synthesizes external network environmental dynamics with internal network requirements, accompanied by the first documented implementation case study based on this architectural framework. Second, it provides empirical validation through a RAN LA Agent functioning at the network element level. This case study represents pioneering real-world demonstrations of AN Agent capabilities, quantitatively validating L4 autonomy through network performance enhancements.
II From AI Agent to AN Agent
II-A Current AI Agent Architectures
Contemporary AI agents achieve environmental interaction through world models that explicitly represent environmental dynamics as learnable transition functions or implicitly encode them within latent neural representations. LeCun’s architecture [3] exemplifies explicit approaches with modular predictive models for real-time control, while implicit paradigms like DayDreamer [4] encode dynamics in latent structures for sample-efficient robot learning. The A3T framework [5] advances this paradigm through autonomous trajectory synthesis, achieving 96% success in embodied tasks without manual intervention.
Modern reasoning agents increasingly rely on LLMs as cognitive cores. The hybrid LAW framework [6] integrates symbolic modules for logical consistency in precision tasks, while ReAct [7] dynamically interleaves cognition with environmental interaction to mitigate hallucinations. Self-Consistency [8] enhances reliability through diverse reasoning path sampling, boosting accuracy by +17.9% on benchmarks. Embodied Reasoning [9] fuses multimodal inputs with actuation for physical adaptation, while Social Reasoning [10] employs graph embeddings for multi-agent coordination. These reveal fundamental trade-offs: symbolic methods ensure precision but constrain flexibility, while pure LLM approaches prioritize adaptability with hallucination risks.
Self-evolution architectures enable autonomous optimization of cognitive structures (e.g., workflows, logic modules) through closed-loop mechanisms, reducing human intervention while improving adaptability. Hu et al. [11] pioneer this with meta-agents programming novel agents in Turing-complete code space. Liu et al. [12] formalize multi-agent collaboration as computational graphs with dynamic pruning. Liu et al. [13] further enhance edge LLM agents via cloud-edge collaboration, where a cloud-hosted LLM and edge agent co-optimize through a lightweight adapter.
While sharing foundational perception-feedback mechanisms, these paradigms reflect distinct evolutionary pathways: computational efficiency (World Models), cognitive integrity (Reasoning Systems), and recursive metamorphosis (Self-Evolution). Future progress requires cross-paradigm synthesis and ethical resolution as systems evolve into sociotechnical partners.
II-B AN Agent Reference Architecture
Figure 1: Sifakis’s Agent architecture with proactive and reactive behaviors.
Joseph Sifakis and collaborators have established a reference architecture [14] (shown in Fig. 1) for autonomous agents that intrinsically supports both proactive and reactive behavioral modalities through integrated cognitive mechanisms. This framework adheres to a fundamental systems engineering paradigm by generalizing existing agent solutions—characterizing autonomy through the composition of mutually independent, mathematically defined cognitive functions, independent of implementation specifics. The architecture’s behavioral completeness captures the essential aspects of human cognition, particularly Kahneman’s dual-process theory [15], while maintaining sufficient generality to model an agent as an entity continuously interacting with its internal state and external environment (collectively the “world”). Here, behavioral competence emerges from two interconnected systems coordinated around a central Long-Term Memory repository: the reactive subsystem generates rapid, Kahneman’s System 1-like responses to environmental perturbations, while the proactive subsystem engages in System 2-style deliberative analysis of internal state variations.
Reactive functionality, extensively documented in domains such as autonomous robotics, operates through sensory input interpretation semantically enriched by contextual knowledge from long-term memory. For instance, when an autonomous vehicle’s perception system identifies a stop sign through visual sensors, this raw sensory input activates associated procedural knowledge (“deceleration imperative”) not intrinsically embedded in the pixel data, demonstrating how contextual augmentation transforms environmental sensing into actionable intelligence. The proactive system architecturally integrates self-awareness and choice-making subsystems that operate in concert. The self-awareness module continuously monitors the agent’s operational state against predefined existential objectives—such as maintaining service availability thresholds in telecommunications contexts—triggering corrective intentionality when deviations exceed tolerance boundaries. These intentions are then mapped to corresponding meta-goals by the self-awareness module. The choice-making module receives these meta-goals and selects between possible goal alternatives through cost-benefit analysis of goals satisfying the identified need.
For operational contexts requiring human supervision or multi-agent collaboration, the architecture incorporates specialized interface extensions. Human-Agent Interaction modules process operator directives and contextual network data through structured dialogue protocols, while Agent-Agent Interaction modules implement standardized coordination mechanisms for collective problem-solving. These extensions significantly enhance the framework’s applicability in complex sociotechnical environments where negotiation and cooperative task execution are essential for mission success.
III Building Autonomous Network Agents: Implementation Framework and Technical Realization
III-A Executable Runtime Instantiation
The architecture of autonomous agents requires a clear separation between functional modules and runtime orchestration. Modules serve as deterministic input-output units—such as perception processors transforming sensory data or planners generating action sequences. Runtime systems, however, act as meta-coordination layers that dynamically sequence modules via state-machine logic. This decoupling allows flexible reconfiguration of the same modules for different domains. As the agent’s cognitive core, runtime systems must interpret heterogeneous inputs (e.g., human directives or environmental telemetry), maintain contextual awareness of available capabilities (including function invocation and knowledge retrieval), and strategically navigate multidimensional solution spaces. These spaces are conceptualized as topological structures linking problem states to resolution pathways, with execution plans representing optimized trajectories through this landscape.
Figure 2: A flat view of an agent.
At the architectural apex resides the Workflow Coordinator Runtime in our implementation for AN, functioning as the supreme coordination authority that dynamically allocates computational resources and governs activation sequences across specialized subsystems. This coordinator maintains a holistic state representation while managing (1) human interfaces for operator interaction, (2) analytical tools including optimization validators and formal verification engines, (3) observation pipelines processing real-time network sensory information, (4) world knowledge systems, (5) external knowledge bases, and domain-specific repositories, and (6) LLMs. This architectural nexus establishes the functional topology shown in Fig. 2, where the coordinator occupies the central hub position of subsystems.
Figure 3: Architectural runtime workflow diagram illustrating the coordination between reactive (processes 1–7) and proactive (processes a–i) runtime systems.
The Workflow Coordinator Runtime can invoke two runtimes: Reactive Behavior Runtime and Proactive Behavior Runtime. The Reactive Behavior Runtime (shown in Fig. 3 processes 1–7) executes time-sensitive adaptations through a structured sensory-processing pipeline. When environmental stimuli are detected, the perception module generates enriched contextual interpretations that update long-term memory. This triggers iterative knowledge retrieval cycles that progressively refine contextual understanding until sufficient information is gathered to construct a predictive model. Using this model, the system generates candidate goals that undergo multi-criteria evaluation. The optimal goal selection then activates constraint-based planners to synthesize executable actions. Proposed actions are validated by the predictive model, forming a closed-loop system. This feedback mechanism continuously adjusts goals based on projected outcomes before finalizing execution plans.
The Proactive Behavior Runtime (shown in Fig. 3 processes a–i) initiates operations through continuous self-monitoring mechanisms that compare current states against predefined operational targets. When deviations exceed critical thresholds, intent generation modules synthesize operational needs by combining intrinsic objectives with externally retrieved constraints from long-term memory. These needs materialize as meta-goals (higher-order objectives), which undergo feasibility assessment against system constraints. The meta-goal is then translated into technical objectives, further justified by feasibility constraints, and the optimal goal is selected via cost-benefit analysis. The chosen goal is forwarded for management and planning, leveraging the same predictive modeling and planning infrastructure used by the reactive runtime. This architectural consistency ensures unified decision semantics across operational modes while supporting complex goal hierarchies.
III-B Technology Selection for Functional Modules
The heterogeneous nature of network knowledge—spanning static protocols, dynamic topologies, and experiential patterns—necessitates a tiered storage approach, combining graph databases (e.g., Neo4j) for structured ontologies (3rd Generation Partnership Project (3GPP) standards, entity relationships) and vector databases (e.g., FAISS) for dynamic operational embeddings. This dual-structured approach balances semantic precision and adaptive scalability, where graph models preserve logical hierarchies while vector spaces enable similarity-based retrieval of evolving patterns. Retrieval operations merge contextual pattern matching—implemented through Retrieval-Augmented Generation (RAG) pipelines that query embeddings for analogous operational states—with symbolic reasoning (Web Ontology Language (OWL)-based inference over ontologies), converging through semantic alignment layers that resolve conflicts via multi-head attention. The system thus maintains logical consistency while adapting to contextual dynamics, synthesizing actionable insights from both empirical precedents and deductive constraints.
The Situation Awareness module confronts real-time sensory noise and predictive uncertainty through a layered analytical pipeline. In general, it contains multimodal data such as time-series measurements, images, audio, or video. In this article, we will focus solely on processing time-series data. To mitigate transient artifacts, signal conditioning employs sliding-window averaging for baseline normalization, Kalman filters for noise-correlated sensor fusion, wavelet thresholding for multiscale transient suppression, median filtering for impulse/spike removal, and robust regression techniques for outlier-resilient feature extraction. These jointly establish stable input baselines for multi-tiered Long Short-Term Memory (LSTM) networks to predict future trends. This transforms denoised inputs into structured JSON summaries encapsulating trend trajectories, vulnerability diagnostics and rule-based goal suggestions.
The Self-Awareness module initiates cognitive processes through three distinct activation pathways: direct human directives via natural language interfaces, periodic state reviews governed by temporal scheduling logic, and event-driven triggers responding to critical network anomalies. Upon activation, intent generation employs iterative knowledge retrieval cycles from long-term memory repositories, where contextualized historical states are fused with current operational metrics. This synthesized context drives a few-shot LLM to generate purpose-aligned intents, constrained by domain-specific templates that enforce syntactic and semantic consistency. Subsequently, meta-goal synthesis extends this reasoning paradigm by further refining the validated intents through additional knowledge retrieval iterations. Leveraging the same LLM framework with mission-adapted prompt engineering, the module produces abstract yet actionable meta-goals that encapsulate strategic objectives while remaining decoupled from implementation specifics.
The Choice Making module converts strategic meta-goals into executable technical objectives using a hierarchical decision process. A lightweight multilayer perceptron (MLP) first generates probabilistically ranked candidate goals by processing encoded state vectors—combining real-time telemetry, historical patterns, and meta-goal constraints—while maintaining sub-millisecond latency for autonomous operation. Optimal goal selection then leverages deep reinforcement learning (DRL), where dynamically weighted KPI combinations parameterize reward functions to resolve multi-objective trade-offs. This model-free approach inherently adapts to dynamic environments through learned value functions, balancing competing operational priorities efficiently.
The Decision Making module synthesizes executable actions by integrating reactive and proactive behaviors. The goal management process resolves semantic conflicts using telecom-ontology enhanced LLMs through contextual reasoning of network dynamics, while rule engines ensure deterministic adherence to protocol specifications through temporal logic verification. The resulting reconciled goal drives Monte Carlo Tree Search (MCTS) planning, where stochastic simulations iteratively select optimal branches while maintaining formal safety guarantees through constraint-aware backpropagation.
IV Case Study: Radio Access Network (RAN) Link Adaptation (LA) Agent
The performance optimization of Radio Access Networks (RAN) constitutes a fundamental challenge in achieving extreme performance metrics (e.g., ultra-reliable low-latency communication (URLLC) and enhanced mobile broadband (eMBB)) in 5G/6G networks. RAN Link Adaptation (LA) technology, due to its sensitivity to dynamic air interface conditions and millisecond-level real-time decision requirements, serves as an ideal validation case for the AN Agent architecture. LA’s primary mechanism involves dynamically selecting the Modulation and Coding Scheme (MCS) - a fundamental control knob that dictates both the throughput (TPT) and reliability on the wireless link. In 5G and future networks, LA must dynamically adjust MCS in complex interference environments (e.g., nonlinear power amplifier effects, inter-cell interference), diverse service requirements (TPT, reliability, latency), and multi-dimensional parameter optimization scenarios to maximize spectral efficiency while ensuring reliability. However, traditional standardized approaches relying on post-facto feedback adjustment mechanisms are too slow to address these challenges. This case study introduces the LA Agent to validate its capability for sub-10 ms real-time closed-loop control at the network element level. By deploying intelligent Agents with local decision-making capabilities, the system continuously monitors key RAN indicators, autonomously analyzes potential air interface bottlenecks, and adjusts LA strategies accordingly.
IV-A Problem Modeling
This case study focuses on the MCS optimization problem in the RAN architecture: selecting the optimal MCS for each transport block (TB) in unknown time-varying channels to maximize TPT while satisfying Block Error Rate (BLER) target constraints. The problem is inherently a real-time sequential decision-making challenge in high-dimensional stochastic environments—for each TB transmission, the base station must predict the BLER induced by candidate MCS under current channel conditions based on limited and delayed channel state information (e.g., Channel Quality Indicator (CQI) derived from User Equipment (UE) measurements). The system must then control BLER to meet both the target threshold and TPT maximization objectives.
Traditional solutions employ the Outer Loop Link Adaptation (OLLA) mechanism: predefined Signal-to-Interference-plus-Noise Ratio (SINR)-BLER mapping tables based on ideal channel models, with dynamic adjustment of SINR offsets via ACK/NACK feedback. However, this approach suffers from three critical limitations: (1) OLLA’s offset adjustment relies on extensive historical feedback (typically requiring dozens of TB cycles), leading to significant convergence delays in low-traffic or high-mobility scenarios, causing prolonged suboptimal MCS selection; (2) predefined SINR-BLER models cannot accurately adapt to real-world nonlinear impairments (e.g., power amplifier distortion, phase noise, dynamic interference), particularly in millimeter-wave multi-beam scheduling scenarios, where model mismatch results in BLER control deviations; and (3) uniform offset strategies fail to accommodate differentiated service requirements (e.g., coexisting eMBB and URLLC services in a single cell requiring BLER targets of $10^{-1}$ and $10^{-3}$, respectively).
IV-B Engineering Details
To address these challenges, we developed a RAN LA Agent based on the reference architecture and our implementation shown in Figs. 13. The design implementation is as follows:
The Long-Term Memory module: employs FAISS for vector storage of real-time channel states (SINR/BLER/TPT) and MCS histories, while Neo4j stores structured 3GPP MCS properties (e.g., QPSK/64-QAM spectral efficiency). Hard-coded rules enforce reward mode switching of eMBB/URLLC via BLER thresholds. Knowledge retrieval implements RAG for optimal MCS recommendation from historical analogs, while a lightweight rule engine executes threshold-based reasoning for protocol compliance.
Situation Awareness Module: sampled data undergo a sliding-window averaging for ACK/NACK/TPT smoothing and Kalman-filtered SINR estimation. LSTM networks project BLER trends across 5 future Transmission Time Intervals (TTIs) using 100-TTI historical inputs.
Self-Awareness Module: governs network operational modes through dual meta-goal formulations: URLLC enforces “BLER $0.1\%$ with minimal latency” while eMBB pursues “maximum TPT under BLER $10\%$ constraints”. These abstract directives emerge from either event-triggered threshold violations or human natural language instructions processed via few-shot LLM interpretation.
Choice Making Module: generates candidate technical goals through a lightweight MLP trained on historical link adaptation decisions, demonstrating distinct meta-goal decomposition patterns across operational modes. In URLLC scenarios, meta-goal realization involves two kinds of goals: MCS modulation order adjustment for reliability optimization and Multiple-Input Multiple-Output (MIMO) transmission rank modification for latency minimization, both dynamically calibrated to current channel conditions. For eMBB operation, the system concurrently pursues three kinds of goals: MCS modulation adaptation, coding rate adjustment, and MIMO rank optimization to exploit spatial multiplexing gain. All proposed actions undergo Deep Q-Network (DQN)-based evaluation with dynamically weighted reward functions that prioritize either BLER reliability or TPT according to active service requirements as determined by the meta-goal. The final selection employs a greedy strategy that chooses the highest-scoring goal based on the weighted evaluation metrics.
Decision Making Module: The module enforces protocol compliance through rule-based priority constraints (e.g., mandatory MCS reduction when BLER $0.1\%$), while directly executing DQN-selected MCS commands to base stations without intermediate planning.
The architecture efficiently coordinates 2-3 core objectives (BLER, TPT, latency) in our RAN scenarios. Complexity is actively managed through two complementary strategies: rule-based pruning eliminates logically conflicting goals (e.g., disabling TPT maximization during URLLC operation), while dynamic reward functions adaptively weight competing KPIs based on active service priorities (URLLC prefers low BLER; eMBB prefers high TPT). Three core challenges are addressed in predictive modeling: (1) goal conflicts, which are mitigated through hierarchical prioritization and DQN penalties for goals that contradict active operational modes; (2) real-time constraints, handled via LSTM-based short-term BLER forecasting; and (3) multi-objective optimization, achieved through service-specific reward functions.
Our current implementation establishes a functional autonomous decision-making pipeline, integrating real-time MCS adaptation (DQN with LLM-adjusted rewards), LSTM-based channel prediction, and RAG-enhanced knowledge retrieval. The system achieves lightweight cognitive capabilities through MLP-based goal proposals and rule-constrained conflict resolution while laying foundations for advanced autonomy via LLM-mediated intent generation. Future extensions will: (1) develop MCTS for multi-step action planning under 3GPP timing constraints, (2) implement telecom-ontology-guided LLM reasoning for complex conflict resolution, (3) expand multimodal signal processing, and (4) integrate OWL-based reasoning for multi-cell interference modeling.
IV-C Experiment Environment and Results
(a) Control computer, Remote Radio Units (RRUs), and Test UEs.
(b) Baseband Unit (BBU).
Figure 4: Test site environment. Lab wireless setup includes: 1x BBU, 2x RRUs, 2x test UEs, and 1x measurement/control host.
The experimental testbed configuration is illustrated in Fig. 4, featuring a control workstation running EaglePro 6.4 software for remote terminal management and comprehensive data acquisition of signaling traces and performance metrics. Two commercial 5G smartphones with distinct chipset architectures were deployed as test terminals to evaluate cross-platform compatibility: a Huawei Mate 40 Pro with software version 2.0.0.990 and a VIVO IQOO 11 running software version PD2243B_A_13.0.11.11.W10.V [^0] L1. The radio access network architecture incorporated two Remote Radio Units (RRUs) connected via fiber-optic fronthaul to an ISAC2 Baseband Unit (BBU), which was equipped with an Intel(R) Xeon(R) Silver 4416+ CPU operating at 2.0 GHz, 128 GB of system memory, and an NVIDIA L20 GPU with 48 GB of dedicated video memory.
This open-air testbed operated in the 3GPP N78 band centered at 3.5 GHz with 100 MHz of allocated bandwidth, maintaining controlled channel conditions characterized by Reference Signal Received Power levels of approximately -73 dBm with ±1 dB variation, Reference Signal Received Quality measurements around -10 dB with ±0.5 dB fluctuation, and SINR values near 27 dB within a ±1 dB tolerance window. These carefully maintained parameters established a stable yet dynamically representative radio environment that effectively simulated realistic 5G and emerging 6G network conditions, where advanced adaptive algorithms must compensate for rapid channel variations and interference patterns.
A rigorous comparative methodology was employed to evaluate the LA Agent against the conventional 3GPP downlink link adaptation OLLA algorithm. Data acquisition occurred via drive-test software, sampling KPIs at 100 ms intervals over a continuous 5-minute duration. KPIs included downlink Packet Data Convergence Protocol (PDCP) average TPT (measured in Mbps) and Physical Downlink Shared Channel (PDSCH) BLER, with values plotted against sample indices for temporal trend analysis. This high-resolution monitoring allowed for granular assessment of transient behaviors and steady-state performance across distinct operational objectives. The evaluation focused on two critical test cases derived from practical network demands: first, maximizing spectral efficiency to optimize TPT under typical enhanced eMBB conditions; and second, maximizing the assurance of PDCP throughput while prioritizing low BLER for URLLC scenarios, thereby testing the agent’s capability to enforce stringent reliability constraints. The results are shown in Fig 5.
Figure 5: Comparative performance evaluation of LA Agent versus OLLA algorithm.
The comparative evaluation of Test Case 1 reveals critical temporal dynamics in the LA Agent’s environmental adaptation capabilities. As shown in the experimental timeline, the agent demonstrated significantly accelerated responsiveness to fluctuating channel conditions compared to conventional 3GPP OLLA mechanisms. During initial high-quality air-interface phases (Sample #30-100), the agent rapidly exploited favorable conditions by elevating MCS levels, achieving improvement in downlink PDCP throughput over the baseline OLLA approach. This responsiveness stems directly from the agent’s integrated reactive behavior runtime, whose sub-10 ms perception-action cycle enables near-instantaneous capitalisation on transient spectral opportunities—a capability unattainable by feedback-delayed OLLA systems requiring dozens of Transport Block (TB) cycles for offset calibration.
Crucially, when channel quality deteriorated abruptly (Sample #100-200), the agent proactively reduced MCS allocation within a single decision cycle to maintain BLER below the eMBB operational threshold. This contrasts sharply with OLLA’s delayed response, which exhibited persistent TPT overshoot and momentary BLER violations during transition periods. The dual-mode adaptability—aggressive spectral exploitation during optimal conditions and conservative reliability enforcement during degradation—validates the architectural synergy between the reactive and proactive runtimes. Quantitative results underscore our advantage, with the LA Agent achieving 504 Mbps average TPT versus OLLA’s 476 Mbps - a statistically significant 6% improvement in spectral efficiency.
For Test Case 2, targeting high-reliability URLLC scenarios with maximized TPT assurance, the LA Agent exhibited exceptional BLER control capabilities. While the conventional OLLA mechanism achieved an average BLER of 0.7%, the LA Agent further reduced this to 0.2% - representing a significant 67% improvement in error rate control. This performance breakthrough stems from the LA Agent’s predictive capability - it proactively adjusts MCS based on anticipated channel conditions, whereas traditional algorithms react only after BLER degradation has already occurred. While the TPT improvement in Case 1 may appear modest compared to the dramatic BLER reduction, this reflects fundamental system characteristics: (1) the baseline TPT was already operating near theoretical maximums because of the relatively stable and slowly-varying channel environment, and (2) the absolute improvement still delivers practically meaningful gains (28Mbps) for real-world deployments.
V Conclusion
This work implements an autonomous network agent architecture through hybrid knowledge representation and coordinated runtime integration. Empirical validation demonstrates 6% gain in TPT and 67% improvement in BLER over OLLA benchmarks while meeting 10 ms latency constraints, overcoming prevailing automation limitations. The demonstrated framework provides a foundation for addressing emerging requirements in 6G networks, while maintaining the rigorous protocol compliance and real-time responsiveness that are essential for practical autonomous operation. This work suggests that continued refinement of hybrid knowledge representation and hierarchical decision-making systems will be crucial for achieving the full potential of autonomous networks in next-generation telecommunications infrastructure.